# PAC-Bayes Analysis for Recalibration in Classification

### v1

@@ A.2.3 Proof of Theorem 2 (Generalization of the ECE under the training dataset)
$=\left|\mathbb{E}_{S_{\mathrm{te}}}\sum_{i=1}^{B}\left|\frac{1}{n}\sum_{(X,Y)\in S_{\mathrm{te}}}\left[(Y-f_{W}(X))\cdot\mathbbm{1}_{f_{W}(X)\in I_{i}}\right]\right|-\sum_{i=1}^{B}\left|\frac{1}{n}\sum_{(X,Y)\in S_{\mathrm{tr}}}(Y-f_{W}(X))\cdot\mathbbm{1}_{f_{W}(X)\in I_{i}}\right|\right|$
@@
where we used the triangle inequality $||a|-|b||\leq|a-b|$ for the first inequality and set $l_{i}(z)=(y-f_{W}(x))\cdot\mathbbm{1}_{f_{W}(x)\in I_{i}}$.
@@
A.3 Proof of Theorem 3 (Estimation bias under the training dataset)
First, we present the complete statement again;
Theorem 7.
Assume that $n_{\mathrm{te}}=n_{\mathrm{tr}}=n$. Under Assumption 1, for any fixed prior $\pi$, which is independent of $S_{\mathrm{tr}}$, and posterior distribution $\rho$ over $W$ and any positive constant $\varepsilon\in[0,1]$ and $\lambda>0$, with probability $1-\varepsilon$ with respect to $S_{\mathrm{tr}}$, we have
$\mathbb{E}_{\rho}\mathrm{Bias}(f_{w},S_{\mathrm{tr}},\mathrm{CE})\leq\frac{1+L}{B}+\frac{\mathrm{KL}(\rho\|\pi)+B\log 2+\log\frac{1}{\varepsilon}+\frac{2\lambda^{2}}{n}}{\lambda},$ (53)
$\mathbb{E}_{\rho}\mathrm{Bias}(f_{w},S_{\mathrm{tr}},\mathrm{TCE})\leq\frac{1+L}{B}+\frac{\mathrm{KL}(\rho\|\pi)+B\log 2+\log\frac{1}{\varepsilon}+\frac{2\lambda^{2}}{n}}{\lambda}.$ (54)
Proof.
Binary setting
We start from the binary classification setting. As stated in the proof outline, we decompose the estimation bias into two biases as follows:
$|\mathrm{CE}(f_{w})-\mathrm{ECE}(f_{w},S_{\mathrm{tr}})|\!\leq|\mathrm{CE}(f_{w})-\mathrm{CE}(f_{\mathcal{I}})|+|\mathrm{CE}(f_{\mathcal{I}})-\mathrm{ECE}(f_{w},S_{\mathrm{tr}})|,$ (55)
where
In the above, the second term, $\mathbb{E}_{\rho}|\mathrm{CE}(f_{\mathcal{I}})-\mathrm{ECE}(f_{w},S_{\mathrm{tr}})|$ has already been proved in Theorem 5. So we only need to bound the first term. As for the first term, we show that
holds. First, we can show that
$\mathrm{CE}(f_{\mathcal{I}})\leq\mathrm{CE}(f_{w}),$ (58)
$\quad+\mathbb{E}[f_{w}(X)|f_{w}(X)\in I_{i}]-f_{w}(X)||f_{w}(X)\in I_{i}]$
Then we can show that
where we used the Lipschitz continuity of the function and the fact that we split the function with equal width $1/B$. Then using Eq. (55) and above result, we have
and use Theorem 5 for the second term, this concludes the proof for the binary setting. ∎
Proof.
Multiclass setting
Next, for the multiclass setting, similarly to the binary case, we introduce the conditional expectation defined as;
As for the multiclass setting, similarly to the binary case, we have
$|\mathrm{TCE}(f_{w})-\mathrm{TCE}(f^{C}_{\mathcal{I}})|\leq\mathbb{E}||\mathbb{E}[Y|f_{w}(X)_{C}]-\mathbb{E}[Y|f_{\mathcal{I}}^{C}(X)]|+\mathbb{E}|f_{w}(X)_{C}-f_{\mathcal{I}}^{C}(X)|$ (64)
by the almost identical proof. We then upper bound them
$|\mathrm{TCE}(f_{w})-\mathrm{TCE}(f^{C}_{\mathcal{I}})|\leq\mathbb{E}||\mathbb{E}[Y|f_{w}(X)_{C}]-\mathbb{E}[Y|f_{\mathcal{I}}^{C}(X)]|+\mathbb{E}|f_{w}(X)_{C}-f_{\mathcal{I}}^{C}(X)|\leq\frac{L}{B}+\frac{1}{B}$ (65)
$\mathbb{E}_{\rho}|\mathrm{TCE}(f_{w})-\mathrm{ECE}(f_{w},S_{\mathrm{tr}})|\!$
Then using Theorem 5 for the second term, this concludes the proof. ∎
A.4 Limitation under the Hölder continuity
Here we discuss the assumption of Lipschitz continuity. It is known that by assuming the higher order smoothness, such as the $\beta$-Hölder continuity, the bias of the nonparametric estimation decreases [30], particularly, the lower bound is $\mathcal{O}(n^{-\frac{\beta}{2\beta+1}})$.
@@ A.4 Limitation under the Hölder continuity
A.5 Proof of Theorem 4 (Estimation bias for the multiclass setting)
The upper bound can be derived similarly to the binary classification setting in Appendix A.3; 1) we first transform the $ECE_{K}$ and $CE_{K}$, 2) decompose the estimation bias to the generalization and binning bias, 3) bound those two terms using the generalization error analysis and the binning definitions.
@@ A.5 Proof of Theorem 4 (Estimation bias for the multiclass setting)
We first derive the ECE transformation; for the $ECE_{K}$ under the training dataset, we have
$\mathrm{ECE}_{K}(f_{w},S_{\mathrm{tr}})=\sum_{i=1}^{B}\frac{\#B_{i}}{n}\|\bar{f}_{i,S_{\mathrm{tr}}}-\bar{y}_{i,S_{\mathrm{tr}}}\|=\sum_{i=1}^{B}\|\mathbb{E}_{(X,\hat{Y})\sim\widehat{S}_{\mathrm{tr}}}(\hat{Y}-f_{w}(X))\cdot\mathbbm{1}_{f_{w}(x_{m})\in I_{i}}\|.$ (67)
@@
$\mathrm{CE}_{K}(f_{\mathcal{I}})=\sum_{i=1}^{B}\|\mathbb{E}_{(X,\hat{Y})\sim\mathcal{D}}(\hat{Y}-f_{w}(X))\cdot\mathbbm{1}_{f_{w}(X)\in I_{i}}\|.$ (69)
@@
Assume that $n_{\mathrm{te}}=n_{\mathrm{tr}}=n$ and $f_{w}(X)$ has the probability density over $[0,1]^{K}$. Then for any fixed prior $\pi$, which is independent of $S_{\mathrm{tr}}$, and posterior distribution $\rho$ over $W$ and any positive constant $\varepsilon\in[0,1]$ and $\lambda>0$, with probability $1-\varepsilon$ with respect to $S_{\mathrm{tr}}$, we have
@@
The proof strategy is almost identical to Appendix A.2.2.
@@
$\mathrm{CE}_{K}(f_{\mathcal{I}})$ $=\sum_{i=1}^{B}\|\mathbb{E}_{(X,\hat{Y})\sim\mathcal{D}}(\hat{Y}-f_{w}(X))\cdot\mathbbm{1}_{f_{w}(X)\in I_{i}}\|$ (72)
@@
$\mathrm{ECE}_{K}(f_{w},S_{\mathrm{tr}})$ $=\sum_{i=1}^{B}\|\mathbb{E}_{(X,\hat{Y})\sim\widehat{S}_{\mathrm{tr}}}(\hat{Y}-f_{w}(X))\cdot\mathbbm{1}_{f_{w}(x_{m})\in I_{i}}\|$ (73)
@@
$=\left|\sum_{i=1}^{B}\|\mathbb{E}_{(X,\hat{Y})\sim\mathcal{D}}(\hat{Y}-f_{w}(X))\cdot\mathbbm{1}_{f_{w}(X)\in I_{i}}\|-\sum_{i=1}^{B}\|\mathbb{E}_{(X,\hat{Y})\sim\widehat{S}_{\mathrm{tr}}}(\hat{Y}-f_{w}(X))\cdot\mathbbm{1}_{f_{w}(x_{m})\in I_{i}}\|\right|$
$\leq\sum_{i=1}^{B}\left\|\mathbb{E}_{(X,\hat{Y})\sim\mathcal{D}}l_{i}(Z)-\frac{1}{n}\sum_{Z\in S_{\mathrm{tr}}}l_{i}(Z)\right\|,$ (74)
where we used the triangle inequality $|\|a\|-\|b\||\leq\|a-b\|$ for the first inequality and set $l_{i}(z)=(\hat{y}-f_{W}(x))\cdot\mathbbm{1}_{f_{W}(x)\in I_{i}}$. We then evaluate the exponential moment as follows;
$\mathbb{E}_{S_{\mathrm{tr}},\pi}e^{t|\mathrm{CE}_{K}(f_{\mathcal{I}})-\mathrm{ECE}_{K}(f_{w},S_{\mathrm{tr}})|}$ $\leq\mathbb{E}_{\pi}\mathbb{E}_{S_{\mathrm{tr}}}\mathbb{E}_{S_{\mathrm{te}}}e^{t\sum_{i=1}^{B}\left\|\mathbb{E}_{(X,\hat{Y})\sim\mathcal{D}}l_{i}(Z)-\frac{1}{n}\sum_{Z\in S_{\mathrm{tr}}}l_{i}(Z)\right\|}$
$\leq\mathbb{E}_{\pi}\mathbb{E}_{S_{\mathrm{tr}}}\mathbb{E}_{S_{\mathrm{te}}}e^{t\sum_{i=1}^{B}\sum_{k=1}^{K}|\mathbb{E}_{(X,\hat{Y})\sim\mathcal{D}}l_{i}(Z)_{k}-\frac{1}{n}\sum_{Z\in S_{\mathrm{tr}}}l_{i}(Z)_{k}|}$ (75)
@@
$=\sup_{z^{\prime}_{m},\tilde{z}_{m^{\prime}}\in\mathcal{Z}}\sum_{k=1}^{K}\frac{t(-1)^{v_{1}}}{n}\cdot\Big{(}-\big{(}(\hat{y}^{\prime}_{m^{\prime},k}-f_{W}(x^{\prime}_{m^{\prime}})_{k})\cdot\mathbbm{1}_{f_{W}(x^{\prime}_{m^{\prime}})\in I_{1}}\big{)}+\big{(}(\hat{\tilde{y}}^{\prime}_{m^{\prime},k}-f_{W}(\tilde{x}^{\prime}_{m^{\prime}})_{k})\cdot\mathbbm{1}_{f_{W}(\tilde{x}^{\prime}_{m^{\prime}})\in I_{1}}\big{)}\Big{)}+$
@@
$+\frac{t(-1)^{v_{B}}}{n}\cdot\Big{(}-\big{(}(\hat{y}^{\prime}_{m^{\prime},k}-f_{W}(x^{\prime}_{m^{\prime}})_{k})\cdot\mathbbm{1}_{f_{W}(x^{\prime}_{m^{\prime}})\in I_{B}}\big{)}+\big{(}(\hat{\tilde{y}}_{m^{\prime},k}-f_{W}(\tilde{x}_{m^{\prime}})_{k})\cdot\mathbbm{1}_{f_{W}(\tilde{x}_{m^{\prime}})\in I_{B}}\big{)}\Big{)}$ (78)
@@
We then repeat the Chernoff-bounding technique of the proof of the PAC-Bayes bound derivation in Appendix A.2.2, and substitute $f(t,n)$ into Eq. (40), we have the following PAC-Bayes bound; for any prior $\pi$ and posterior distribution $\rho$ over $W$ and any positive constant $\varepsilon\in[0,1]$ and $\lambda>0$, with probability $1-\varepsilon$ with respect to $S_{\mathrm{tr}}$, we have
@@
$|\mathrm{CE}_{K}(f_{w})-\mathrm{CE}_{K}(f_{\mathcal{I}})|\leq\mathbb{E}\|\mathbb{E}[\hat{Y}|f_{w}(X)]-\mathbb{E}[\hat{Y}|f_{\mathcal{I}}(X)]\|+\mathbb{E}\|f_{w}(X)-f_{\mathcal{I}}(X)\|.$ (83)
@@
Then we use the fact that $B=(B^{\prime})^{L}$, we have that
@@
$|\mathrm{CE}_{K}(f_{w})-\mathrm{CE}_{K}(f_{\mathcal{I}})|$ $\leq\mathbb{E}||\mathbb{E}[\hat{Y}|f_{w}(X)]-\mathbb{E}[\hat{Y}|f_{\mathcal{I}}(X)]|+\mathbb{E}|f_{w}(X)-f_{\mathcal{I}}(X)|$
@@ Conclusion
Finally, taking the expectation by $\rho$ in Eq. (70), we upper bound the first and second term using Eq. (A.5) and Eq. (82), we obtain the theorem.
A.6 Discussion about the assumption of Theorem 4
In Theorem 4, we assumed that $f_{w}(x)$ has the probability density in $[0,1]^{K}$. As discussed in Appendix A.2.2, in the binary classification and TCE analysis, this assumption is not required, and if we assume this, our bound results in the improvement of the coefficient, not affect the order of the bound.

### v2

@@ B.6 Generalization of the ECE under the training dataset
$=\left|\mathbb{E}_{S_{\mathrm{te}}}\sum_{i=1}^{B}\left|\frac{1}{n}\sum_{(X,Y)\in S_{\mathrm{te}}}\left[(Y-f_{W}(X))\cdot\mathbbm{1}_{f_{w}(X)_{C}\in I_{i}}\right]\right|-\sum_{i=1}^{B}\left|\frac{1}{n}\sum_{(X,Y)\in S_{\mathrm{tr}}}(Y-f_{W}(X))\cdot\mathbbm{1}_{f_{w}(X)_{C}\in I_{i}}\right|\right|$
@@
where we used the triangle inequality $||a|-|b||\leq|a-b|$ for the first inequality and set $l_{i}(z)=(y-f_{W}(x))\cdot\mathbbm{1}_{f_{w}(X)_{C}\in I_{i}}$.
@@
B.7 Limitation under the Hölder continuity
Here we discuss the assumption of Lipschitz continuity. It is known that by assuming the higher order smoothness, such as the $\beta$-Hölder continuity, the bias of the nonparametric estimation decreases [Tsybakov, 2008], particularly, the lower bound is $\mathcal{O}(n^{-\frac{\beta}{2\beta+1}})$.
@@ B.7 Limitation under the Hölder continuity
B.8 Proof of Theorem 3 (Bias for the multiclass setting)
The upper bound can be derived similarly to the binary classification setting in Appendix B.3; 1) we first transform the $ECE_{K}$ and $\mathrm{CE}_{K}$, 2) decompose the estimation bias to the generalization and binning bias, 3) bound those two terms using the generalization error analysis and the binning definitions.
@@ B.8 Proof of Theorem 3 (Bias for the multiclass setting)
Recall the definition
$\mathrm{ECE}_{K}(f_{w},S_{e})\coloneqq\sum_{i=1}^{B}p_{i}\|\bar{f}_{i,S_{e}}-\bar{p}_{i,S_{e}}\|$ (86)
where $|I_{i}|\coloneqq\sum_{m=1}^{n_{e}}\mathbbm{1}_{f_{w}(x_{m})\in I_{i}}$, $\hat{p}_{i}\coloneqq|I_{i}|/n_{e}$, $\bar{f}_{i,S_{e}}\coloneqq\frac{1}{|I_{i}|}\sum_{m=1}^{n_{e}}\mathbbm{1}_{f_{w}(x_{m})\in I_{i}}f_{w}(x_{m})$, and $\bar{p}_{i,S_{e}}\coloneqq\frac{1}{|I_{i}|}\sum_{m=1}^{n_{e}}\mathbbm{1}_{f_{w}(x_{m})\in I_{i}}\mathbf{e}_{y_{m}}$.
Using this definition, we first derive the ECE transformation; for the $ECE_{K}$ under the training dataset, we have
$\mathrm{ECE}_{K}(f_{w},S_{\mathrm{tr}})=\sum_{i=1}^{B}\frac{|I_{i}|}{n}\|\bar{f}_{i,S_{\mathrm{tr}}}-\bar{p}_{i,S_{e}}\|=\sum_{i=1}^{B}\|\mathbb{E}_{(X,Y)\sim S_{\mathrm{tr}}}(\mathbf{e}_{Y}-f_{w}(X))\cdot\mathbbm{1}_{f_{w}(x_{m})\in I_{i}}\|.$ (87)
@@
$\mathrm{CE}_{K}(f_{\mathcal{I}})=\sum_{i=1}^{B}\|\mathbb{E}_{(X,Y)\sim\mathcal{D}}(\mathbf{e}_{Y}-f_{w}(X))\cdot\mathbbm{1}_{f_{w}(X)\in I_{i}}\|.$ (89)
@@
Assume that $f_{w}(X)$ has the probability density over $[0,1]^{K}$. Then for any fixed prior $\pi$, which is independent of $S_{\mathrm{tr}}$, and posterior distribution $\rho$ over $W$ and any positive constant $\varepsilon\in[0,1]$ and $\lambda>0$, with probability $1-\varepsilon$ with respect to $S_{\mathrm{tr}}$, we have
@@
The proof strategy is almost identical to the TCE.
@@
$\mathrm{CE}_{K}(f_{\mathcal{I}})$ $=\sum_{i=1}^{B}\|\mathbb{E}_{(X,Y)\sim\mathcal{D}}(\mathbf{e}_{Y}-f_{w}(X))\cdot\mathbbm{1}_{f_{w}(X)\in I_{i}}\|$ (92)
@@
$\mathrm{ECE}_{K}(f_{w},S_{\mathrm{tr}})$ $=\sum_{i=1}^{B}\|\mathbb{E}_{(X,Y)\sim\widehat{S}_{\mathrm{tr}}}(\mathbf{e}_{Y}-f_{w}(X))\cdot\mathbbm{1}_{f_{w}(x_{m})\in I_{i}}\|$ (93)
@@
$=\left|\sum_{i=1}^{B}\|\mathbb{E}_{(X,Y)\sim\mathcal{D}}(\mathbf{e}_{Y}-f_{w}(X))\cdot\mathbbm{1}_{f_{w}(X)\in I_{i}}\|-\sum_{i=1}^{B}\|\mathbb{E}_{(X,Y)\sim\widehat{S}_{\mathrm{tr}}}(\mathbf{e}_{Y}-f_{w}(X))\cdot\mathbbm{1}_{f_{w}(x_{m})\in I_{i}}\|\right|$
$\leq\sum_{i=1}^{B}\left\|\mathbb{E}_{(X,Y)\sim\mathcal{D}}l_{i}(Z)-\frac{1}{n}\sum_{Z\in S_{\mathrm{tr}}}l_{i}(Z)\right\|,$ (94)
where we used the triangle inequality $|\|a\|-\|b\||\leq\|a-b\|$ for the first inequality and set $l_{i}(z)=(\mathbf{e}_{Y}-f_{W}(x))\cdot\mathbbm{1}_{f_{w}(X)\in I_{i}}$. We then evaluate the exponential moment as follows;
$\mathbb{E}_{S_{\mathrm{tr}},\pi}e^{t|\mathrm{CE}_{K}(f_{\mathcal{I}})-\mathrm{ECE}_{K}(f_{w},S_{\mathrm{tr}})|}$ $\leq\mathbb{E}_{\pi}\mathbb{E}_{S_{\mathrm{tr}}}\mathbb{E}_{S_{\mathrm{te}}}e^{t\sum_{i=1}^{B}\left\|\mathbb{E}_{(X,Y)\sim\mathcal{D}}l_{i}(Z)-\frac{1}{n}\sum_{Z\in S_{\mathrm{tr}}}l_{i}(Z)\right\|}$
$\leq\mathbb{E}_{\pi}\mathbb{E}_{S_{\mathrm{tr}}}\mathbb{E}_{S_{\mathrm{te}}}e^{t\sum_{i=1}^{B}\sum_{k=1}^{K}|\mathbb{E}_{(X,Y)\sim\mathcal{D}}l_{i}(Z)_{k}-\frac{1}{n}\sum_{Z\in S_{\mathrm{tr}}}l_{i}(Z)_{k}|}$ (95)
@@
$=\sup_{z^{\prime}_{m},\tilde{z}_{m^{\prime}}\in\mathcal{Z}}\sum_{k=1}^{K}\frac{t(-1)^{v_{1}}}{n}\cdot\Big{(}-\big{(}((\mathbf{e}_{Y})^{\prime}_{m^{\prime},k}-f_{W}(x^{\prime}_{m^{\prime}})_{k})\cdot\mathbbm{1}_{f_{W}(x^{\prime}_{m^{\prime}})\in I_{1}}\big{)}+\big{(}(\hat{\tilde{y}}^{\prime}_{m^{\prime},k}-f_{W}(\tilde{x}^{\prime}_{m^{\prime}})_{k})\cdot\mathbbm{1}_{f_{W}(\tilde{x}^{\prime}_{m^{\prime}})\in I_{1}}\big{)}\Big{)}+$
@@
$+\frac{t(-1)^{v_{B}}}{n}\cdot\Big{(}-\big{(}((\mathbf{e}_{Y})^{\prime}_{m^{\prime},k}-f_{W}(x^{\prime}_{m^{\prime}})_{k})\cdot\mathbbm{1}_{f_{W}(x^{\prime}_{m^{\prime}})\in I_{B}}\big{)}+\big{(}(\hat{\tilde{y}}_{m^{\prime},k}-f_{W}(\tilde{x}_{m^{\prime}})_{k})\cdot\mathbbm{1}_{f_{W}(\tilde{x}_{m^{\prime}})\in I_{B}}\big{)}\Big{)}$ (98)
@@
We then repeat the Chernoff-bounding technique of the proof of the PAC-Bayes bound derivation, and substitute $f(t,n)$ into Eq. (68), we have the following PAC-Bayes bound; for any prior $\pi$ and posterior distribution $\rho$ over $W$ and any positive constant $\varepsilon\in[0,1]$ and $\lambda>0$, with probability $1-\varepsilon$ with respect to $S_{\mathrm{tr}}$, we have
@@
$|\mathrm{CE}_{K}(f_{w})-\mathrm{CE}_{K}(f_{\mathcal{I}})|\leq\mathbb{E}\|\mathbb{E}[\mathbf{e}_{Y}|f_{w}(X)]-\mathbb{E}[\mathbf{e}_{Y}|f_{\mathcal{I}}(X)]\|+\mathbb{E}\|f_{w}(X)-f_{\mathcal{I}}(X)\|.$ (103)
@@
Then we use the fact that $B=(B^{\prime})^{K}$, we have that
@@
$|\mathrm{CE}_{K}(f_{w})-\mathrm{CE}_{K}(f_{\mathcal{I}})|$ $\leq\mathbb{E}||\mathbb{E}[\mathbf{e}_{Y}|f_{w}(X)]-\mathbb{E}[\mathbf{e}_{Y}|f_{\mathcal{I}}(X)]|+\mathbb{E}|f_{w}(X)-f_{\mathcal{I}}(X)|$
@@ Conclusion
Finally, taking the expectation by $\rho$ in Eq. (90), we upper bound the first and second term using Eq. (B.8) and Eq. (102), we obtain the theorem.
As for the total bias
$\mathrm{Bias}(f_{w},S_{\mathrm{te}},\mathrm{CE}_{K})\coloneqq|\mathrm{CE}_{K}(f_{w})-\mathrm{ECE}_{K}(f_{w},S_{\mathrm{te}})|$ (107)
We can derive this using the above generalization bound for $\mathbb{E}_{\rho}|CE_{K}(f_{\mathcal{I}})-\mathrm{ECE}_{K}(f_{w},S_{\mathrm{tr}})|$ similarly to the derivation in Appendix B.5. The result is simply dropping the KL term in Eq. (102) and obtain the order $\mathrm{Bias}(f_{w},S_{\mathrm{te}},\mathrm{CE}_{K})=\mathcal{O}(1/n_{\mathrm{te}}^{1/(K+2)})$ under $B=\mathcal{O}(n_{\mathrm{te}}^{\frac{1}{K+2}})$.
B.9 Discussion about the assumption of Theorem 3
In Theorem 3, we assumed that $f_{w}(x)$ has the probability density in $[0,1]^{K}$. In the binary classification and TCE analysis, this assumption is not required, and if we assume this, our bound results in the improvement of the coefficient, not affect the order of the bound.